# Merlin: A Vision Language Foundation Model for 3D Computed Tomography

### v1

@@ 2.3 Zero-shot Cross-Modal Retrieval
Zero-shot cross-modal retrieval evaluates the model’s ability to match a CT image with the corresponding radiology report findings or impressions section and vice versa. In Figure 4b, we plot the distribution of radiology report findings and impressions lengths. We find that 21% of findings sections have lengths that exceed 512 tokens, motivating our choice of the clinical Longformer Li et al. (2022) text encoder (OpenCLIP Cherti et al. (2022) and BioMedCLIP Zhang et al. (2023b) only allow maximum token lengths of 77 and 256, respectively). We find that Merlin significantly (p <<much-less-than$<$ < 0.01) outperforms OpenCLIP and BioMedCLIP on the task of 3D retrieval for retrieving the correct findings out of 64 cases (Image $\rightarrow$ Findings in Figure 4c). We find similar performance for retrieving the correct CT scan out of 64, given a findings section (Findings $\rightarrow$ Image in Figure 4c).
@@ 2.5 Radiology Report Generation
3D semantic segmentation evaluates whether Merlin learns geometric information about various anatomical structures. We find that with 10% of training cases, Merlin performs 11% better than the second best model variation, which is the Merlin architecture with ImageNet I3D initialization (Figure 7b). This demonstrates that Merlin pretraining is particularly beneficial in label scarce scenarios. Although the advantage is reduced with 100% of training labels, Merlin still outperforms the other model variations (Figure 7b). Figure 7c shows that, where performance of baseline models tends to saturate for larger organs (e.g. liver, stomach, and spleen), Merlin performs best relative to baselines on smaller organs or organs with complicated shapes, like the duodenum, small bowel, T12 vertebrae, L1 vertebrae, and gallbladder. Figure 7d presents qualitative segmentation results, where the red arrows indicate differences in the predicted segmentations relative to the ground truth. We observe that the randomly initialized model makes minor mistakes in all three scans, while Merlin makes a mistake in one of the three scans. Top row, center column: the randomly initialized model misses a small part of the caudate liver lobe. Middle row, center and right columns: both Merlin and randomly initialized models miss small portions of the proximal jejunum. Bottom row, center column: the randomly initialized model includes the bottom calyces of the right kidney in the segmentation mask.
@@
Figure 7: 3D semantic segmentation. (a) To adapt Merlin for segmentation, we add a decoder and skip connections between the Merlin encoder and decoder. (b) We compare model variations using average Dice score across 20 organs that appear in abdominal CT. We compare performance of models trained using 100% of training cases and also simulate the data scarce regime with 10% of training cases. (c) We report Dice scores for 20 organs across 4 model variations using 100% of training cases. (d) We qualitatively compare segmentations between the ground truth labels, a model with the Merlin architecture that has been randomly initialized, and Merlin. The red arrows indicate mistakes made by the model relative to the ground truth. Each row is a different patient sampled from the Total Segmentator Wasserthal et al. (2023) test set.
@@ 3 Discussion
In this study, we curate a high-quality clinical dataset comprising of paired CT scans (10,628,509 total 2D images from 25,528 CT scans), EHR data (2,995,293 ICD codes), and radiology reports (10,051,571 total tokens in the findings sections). With this dataset, we train Merlin, a 3D vision-language foundation model for interpreting abdominal CT scans that leverages structured EHR data and unstructured radiology report supervision. Merlin is trained on a single GPU, demonstrating how hospitals and research institutions can pragmatically build their own models, as well as highlighting the opportunity to efficiently train larger models with additional compute and data. Despite the minimal resources used to train Merlin, we demonstrate that Merlin can generalize to 6 types of downstream tasks with 752 individual tasks. Non-adapted tasks include zero-shot findings classification (31 classes), phenotype classification (692 classes), and zero-shot cross-modal retrieval (including both image to findings and image to impressions retrieval). Model adapted tasks include 5-year disease prediction (6 classes), radiology report generation, and 3D semantic segmentation (20 organs). We outperform carefully chosen single-task baseline models across all tasks. We further contextualize our methods with comprehensive ablation studies that demonstrate Merlin’s performance as a function of the underlying model architecture and the design considerations for the use of the underlying data. For training our Merlin foundation model, we demonstrate the benefit of I3D weight initialization, multi-task learning with EHR and reports, and splitting the radiology reports into anatomical sections. Overall, we depict how Merlin may assist in interpretation of abdominal CT scans and mitigate the burden on radiologists, while simultaneously adding value for new biomarker discovery and future disease risk stratification using pre-existing datasets.
@@
Recent concurrent research has started to investigate 3D vision language models for radiology, particularly focusing on chest and head CT scans Hamamci et al. (2024); Yang et al. (2024). However, the extent of experiments and 3D evaluations are limited. These studies lack systematic ablation studies and do not utilize all available clinical data, including EHR. In addition to exploring the use of multiple clinical data types for supervision, we carry out a wide array of experiments and evaluations that contextualize the methodological choices we make in the development of Merlin.
Through the ablation studies, we find that I3D weight initialization Carreira and Zisserman (2017) is helpful for all tasks that we examine with ablation studies. Furthermore, training with both EHR diagnoses and radiology reports is beneficial over training with either alone. The manner of combining these supervision sources is important, with multi-task training using EHR and radiology reports outperforming staged training. Additionally, we find that report splitting, while essential for zero-shot classification, slightly degrades zero-shot retrieval performance. This reinforces that pretraining is most effective when the pretraining data distribution matches the downstream task distribution. Splitting the radiology reports results in text that better aligns with prompts for zero-shot classification, but less closely resembles the radiology report text used for retrieval tasks.
@@
    Increasing dataset size: Expectedly, our initial data scaling results show that using a larger pretraining dataset improves Merlin’s performance (Figure 3b, Figure 4e, Figure 2d). Our data scaling curves across tasks can serve as a useful prospective reference measure in evaluating adequate sizes of training datasets for a requisite task performance. We note that as we add more data, we may need to update our model architecture, which was optimized at our current data scale.
@@
Beyond the strengths of our study, there are also limitations of our model and our work. First, towards our goal of democratization of foundation models, we train and evaluate Merlin using a single GPU. This may limit the generalizability of some of our findings as task performance may increase by simply scaling up compute resources. Second, the limited number of publicly-available baselines for abdominal CT, especially VLM baselines, makes it challenging to understand the relative efficacy of our methods. On the other hand, we believe that Merlin could establish a strong and task-agnostic initial baseline for abdominal CT. Third, further exploration is required to optimize Merlin for the tasks of radiology report generation and 3D segmentation. For radiology report generation, there are numerous parameters that can be tuned, primarily regarding the adapter and LLM. While we devise an initial performant strategy for direct report generation, we leave further optimization of the adapter and LLM to future work. Likewise, adapting Merlin for segmentation requires adding a decoder network. Future work can explore additional architectural parameters and optimisation strategies for decoder training.

### v2

@@ 2.3 Zero-shot Cross-Modal Retrieval
Zero-shot cross-modal retrieval evaluates the model’s ability to match a CT image with the corresponding radiology report findings or impressions section and vice versa. In Figure 4b, we plot the distribution of radiology report findings and impressions lengths. We find that 21% of findings sections have lengths that exceed 512 tokens, motivating our choice of the clinical Longformer Li et al. (2022) text encoder (OpenCLIP Cherti et al. (2022) and BioMedCLIP Zhang et al. (2023b) only allow maximum token lengths of 77 and 256, respectively). We find that Merlin significantly (p <<<< 0.01) outperforms OpenCLIP and BioMedCLIP on the task of 3D retrieval for retrieving the correct findings out of 64 cases (Image $\rightarrow$ Findings in Figure 4c). We find similar performance for retrieving the correct CT scan out of 64, given a findings section (Findings $\rightarrow$ Image in Figure 4c). We also find that on even larger pools of data that Merlin outperforms existing baselines (Supplemental Figure 12).
@@ 2.5 Radiology Report Generation
3D semantic segmentation evaluates whether Merlin learns geometric information about various anatomical structures. We conduct all experiments within the nnUNet framework to ensure a fairer comparison between model architecture and weight initialization by standardizing data-preprocessing, training, and inference. We find that with 10% of training cases, Merlin outperforms nnUNet by 4.7% in average Dice score (Figure 7b). This demonstrates that Merlin pretraining is particularly beneficial in label scarce scenarios. However, the nnUNet model in the 100% training data domain performs slightly better than the Merlin-initialized model, with a Dice score difference of approximately 0.006 (Figure 7b). Among the different initialization strategies for the Resnet152 encoder-decoder framework (ImageNet, random, and Merlin), the performance difference is around 0.01. However, we observe a much larger performance gap, a 6.3% decrease in Dice score, between the fully convolutional segmentation models and the Swin UNETR architecture for this task. Overall, Merlin’s weight initialization outperforms other Resnet-152 initialization strategies in both the 10% and 100% training data settings and nnUNet in the 10% training data setting across all 20 organs.
Across 20 different organs in the test set, we compared the average Dice score of segmentation models (Figure 7c). On a per-organ basis, in the 10% training data domain, Merlin achieves better performance than nnUNet on 12 out of 20 organs, with one organ showing an improvement of over 10% and five organs having a 5% improvement. Notably, Merlin demonstrates a 41.0% improvement in Dice score on the prostate compared to nnUNet. Figure 7d presents qualitative segmentation results, where the red arrows indicate differences in the predicted segmentations relative to the ground truth. We observe that nnUNet makes mistakes in all three slices, while Merlin makes a mistake in two of the three slices. Top row, center column: nnUNet undersegments a small part of the descending colon. Middle row, center and right columns: nnUNet oversegments the urinary bladder and Merlin undersegments the inferior small bowel. Bottom row, center column: both nnUNet and Merlin improperly segment the urinary bladder.
The slightly improved segmentation performance of nnUNet over Merlin in the 100% domain may be attributed to two features. nnUNet’s 3D full-resolution architecture is shallower compared to the 152 layers of Merlin’s Resnet152 (optimized for discriminative tasks), which may reduce potential overfitting. Furthermore, the hyperparameters and pre-processing in the nnUNet framework are specifically tuned for segmentation for a 3D full-resolution architecture, ensuring that the effective receptive field size is appropriately large for the patch size Isensee et al. (2021). This contrasts with the weight-initialized, inflated Resnet152 architecture used in Merlin, which outperforms nnUNet in the 10% training data setting while maintaining state-of-the-art performance across numerous tasks.
@@
Figure 7: 3D semantic segmentation. (a) To adapt Merlin and other architectures for segmentation, we add a decoder and skip connections between the encoder and decoder. We conduct all segmentation experiments within the nnUNet framework. (b) We compare model variations using average Dice score across 20 organs that appear in abdominal CT. We compare performance of models trained using 100% of training cases and also simulate the data scarce regime with 10% of training cases. (c) We report Dice scores for 20 organs across 5 model variations using 10% of training cases. (d) We qualitatively compare segmentations between the ground truth labels, nnUNet, and Merlin with 10% of training cases. The red arrows indicate inconsistencies made by the model relative to the ground truth. The same patient was sampled from the Total Segmentator Wasserthal et al. (2023) test set.
2.7 Alternative Architecture Experiments
Our results from linear probe experiments demonstrate the advantages of Merlin’s 3D vision-language pretraining strategy over 2D finetuned VLMs, 2D-to-3D lifted VLMs, 3D vision-only models, Microsoft’s MedImageInsight model Codella et al. (2024), and Google CT foundation model Yang et al. (2024) (Figure 8). We further show advantages of Merlin pretraining in the data-scarce (10% training data setting) on the findings-based disease classification task (Figure 8d) and on the EHR phenotypes task (Supplemental Figure 11b). These performance gains were also observed in full finetuning evaluations documented in Supplementary Section C.1. Specifically, through our experiments, we observe that: (i) vision-language pretraining outperforms vision-only pretraining, (ii) finetuning improves the performance of 2D and 2D-to-3D VLMs though they underperform relative to Merlin, (iii) Merlin outperforms recent CT embedding foundation models (Figure 8).
2D VLMs: The 2D VLMs demonstrated relatively consistent performance between both finetuning and no finetuning scenarios on the linear probe classification tasks, with finetuned 2D ResnetCLIP performing comparably to the other finetuned models. For example, when trained on 100% of the data for the findings-based disease classification task, ResnetCLIP achieved an F1 score of 0.427 (95% CI [0.381-0.470]), which was similar to OpenCLIP (0.423, 95% CI [0.381-0.463]) and BiomedCLIP (0.424, 95% CI [0.378-0.0.470]). However, 2D VLMs underperformed compared to Merlin, with a 54.7% lower F1 score for findings-based disease classification task and 59.3% lower AUROC on the EHR phenotypes task in the 100% training data setting (Figure 8d-e).
2D-to-3D lifted VLMs: 2D-to-3D lifted VLMs showed consistent performance across tasks and demonstrated performance improvements compared to their 2D VLM counterparts. On the findings-based disease classification task, they achieved a 19.7% increase in F1 score and a 42.9% increase in AUPRC compared to finetuned 2D VLMs in the 100% training data setting (Figure 8d). Similarly, on the EHR phenotype task, 2D-to-3D lifted VLMs outperformed their 2D VLMs baselines, with a 14.2% increase in AUROC and a 47.0% increase in AUPRC (Figure 8e). Despite these improvements, 2D-to-3D lifted VLMs underperformed relative to Merlin. Specifically, Merlin had 32.1% higher F1 scores on the findings-based disease classification task and a 39.8% higher AUROC on the EHR phenotype classification task compared to the 2D-to-3D lifted VLMs (Figure 8d-e).
3D vision-only models: 3D models trained from scratch or with SSL pretraining had a 41.1% decrease in F1 score on the findings-based disease classification task and a 35.8% decrease in AUROC on the EHR phenotypes task compared to Merlin in the 100% training data setting (Figure 8d-e). Among the 3D vision-only baselines, SwinUNETR trailed 3D Resnet SSL, with a 6.3% decrease in F1 score on the findings-based disease classification task and a 18.0% decrease in AUROC on the EHR phenotypes task. The two 3D Resnet models underperformed relative to Merlin, with a 36.9% decrease in F1 score on the findings-based disease classification task and a 28.6% decrease in AUPRC on the EHR phenotypes task (Figure 8d-e). Although 3D vision-only models outperformed the 2D and 2D-to-3D VLMs with a 6.02% increase in F1 score on the findings-based disease classification task and 12.0% increase in AUROC on the EHR phenotypes task, Merlin still achieved the best quantitative performance across all tasks (Figure 8d-e).
CT Embedding Foundation Models: Although MedImageinsight showed a 52.5% increase in AUROC from the 10% to 100% training data setting on the EHR phenotype task (Supplemental figure 11), it still underperformed relative to Merlin with a 33.8% lower F1 on the findings-based disease classification task and 16.6% lower AUROC on the EHR phenotype classification task (Figure 8d-e). Merlin also outperformed the 3D Google CT foundation model, achieving a 5.01% higher F1 and 7.68% higher AUROC on the findings-based disease and EHR phenotype tasks (Figure 8d-e). We therefore find that all alternative architecture baselines underperform relative to Merlin across all label settings and tasks.
Figure 8: Alternative architecture experiments. The baselines, which include three 2D-to-3D lifted vision-language models, five 2D vision-language models, and two 3D vision-only models, are presented to evaluate different training strategies for the AbCT dataset. The baselines serve as points of comparison against Merlin. Below, we illustrate the training procedure for the baseline models. (a) 2D vision-language model (VLM) finetuning engages both the image and text encoders. Specifically, we adapted the image encoder to process randomly selected axial slices from abdominal CT scans and trained the model in a multi-task learning framework, jointly leveraging EHR phenotypes and radiology reports. (b) We finetuned the 2D-to-3D lifted VLM by decomposing the abdominal CT volume into individual slices, which were processed through the image encoder. A neural network slice projector then combined the slice-wise embeddings into a single embedding. (c) Vision-only SSL pretraining using SwinUNETR and Resnet152 entails patchifying a large 3D volume into 3D patches and applying patch-level augmentations, with inpainting, contrastive, and rotational losses. (d) Average F1 (left chart) and AUPRC (right chart) on 10% and 100% pretraining data for the findings-based disease classification task. (e) Average AUROC (left chart) and AUPRC (right chart) across all 1692 phenotypes, the top quartile of 173 phenotypes, and the bottom quartile of 173 phenotypes across several baseline models. All baseline models are trained using the phenotypes in the 100% pretraining dataset. The dashed lines denote random chance performance. Note that Merlin is the best-performing model between these two tasks.
2.8 External Validation Experiments
We draw from a dataset of over 100,000 external CT scans to evaluate Merlin on a total of 44,098 external CT scans, comprising 37,855 abdominal CTs from three external sites and 6,243 chest CTs from site #3. Merlin demonstrates consistent and accurate performance across sites and anatomies, overcoming distributional shifts between training and external testing sets (Supplemental Table  12), and consistently outperforms alternative architecture baselines and chest CT foundation models (Figure  9).
Zero-shot findings classification: We evaluated Merlin on the 30 abdominal findings part of our zero-shot evaluation framework, without any finetuning, using abdominal CT scans for our three external datasets (Figure 9) to assess the robustness of its embeddings to out-of-distribution data sources. For comparison, we include alternative architectures 2D/2D-to-3D BiomedCLIP, OpenCLIP, and ResnetCLIP. Across all sites, Merlin outperformed the second-best alternative architecture baseline by 19.7% on average F1 and demonstrated consistent disease-specific performance across external cohorts (Figure  9a-b). Specifically, at out-of-state site #1, Merlin exceeded 2D-to-3D BiomedCLIP by 34.4%; at in-state site #2, it outperformed 2D-to-3D ResnetCLIP by 15.7%; and at in-state site #3, it maintained an 8.9% advantage over 2D-to-3D ResnetCLIP. These differences largely reflect distribution shifts in patient age, slice thickness, scanner manufacturer, and radiologist reporting patterns (Supplemental Table  12). We attribute Merlin’s improved performance here to its full 3D image encoder pretraining.
Chest CT linear probe evaluation. We evaluate Merlin against recent chest CT FMs Hamamci et al. (2024b); Niu et al. (2025); Pai et al. (2025) and find that, even when training only a linear probe on a frozen Merlin image encoder, Merlin outperforms chest CT FMs by 12.3% on average AUC (Figure 9c-d). Given that Merlin, which was exclusively trained on abdominal CT scans, outperforms other vision-language models trained explicitly on chest CT scans, this demonstrates the generalizability of the Merlin model. Specifically, Merlin surpasses CT-CLIP by 24.7%, M3FM by 14.0%, and achieves equivalent performance to CT-FM. We attribute Merlin’s strong performance on this task to this combination of full-volume 3D processing and diverse multi-modal pretraining.
Figure 9: External validation on abdominal and chest CTs. Merlin’s strong performance across 44,098 abdominal and chest CTs from external sites. (a) Merlin’s performance across internal and three external sites across 30 findings assessed on abdominal CT scans, where Merlin achieved the highest performance across all datasets. (b) Performance breakdown of the second-best model at each site (selected by second-highest average F1) on the same abdominal CT findings. (c) We train a series of embedding models on External Site #3’s chest CT dataset and benchmark them against recent chest CT foundation models. We here present AUC performance comparison for each model across 24 diseases on External Site #3’s dataset. (d) Average AUC performance of Merlin compared to recent Chest CT FMs on External Site #3’s dataset. Across all baselines, Merlin consistently achieves the highest performance. We note that Merlin was trained exclusively on abdominal CT scans, while all other comparison models were trained explicitly on chest CT scans. Despite this training anatomy discordance, Merlin exhibits the strongest performance on Chest CT scans out of all comparison models.
@@ 3 Discussion
In this study, we curate a high-quality clinical dataset comprising of paired CT scans (10,628,509 total 2D images from 25,528 CT scans), EHR data (2,995,293 ICD codes), and radiology reports (10,051,571 total tokens in the findings sections). With this dataset, we train Merlin, a 3D vision-language foundation model for interpreting abdominal CT scans that leverages structured EHR data and unstructured radiology report supervision. Merlin is trained on a single GPU, demonstrating that compute-constrained hospitals and research institutions can build their own models, as well as highlighting the opportunity to train larger models with additional compute and data. We demonstrate that Merlin can generalize to 6 types of downstream tasks with 752 individual tasks. Non-adapted tasks include zero-shot findings classification (31 classes), phenotype classification (692 classes), and zero-shot cross-modal retrieval (including both image to findings and image to impressions retrieval). Model adapted tasks include 5-year disease prediction (6 classes), radiology report generation, and 3D semantic segmentation (20 organs). We outperform carefully chosen single-task baseline models across all tasks. We further contextualize our methods with comprehensive ablation studies that demonstrate Merlin’s performance as a function of the underlying model architecture and the design considerations for the use of the underlying data. For training our Merlin foundation model, we demonstrate the benefit of I3D weight initialization, multi-task learning with EHR and reports, and splitting the radiology reports into anatomical sections. Overall, we depict how Merlin may assist in interpretation of abdominal CT scans and mitigate the burden on radiologists, while simultaneously adding value for new biomarker discovery and future disease risk stratification using pre-existing datasets.
Unlike traditional radiology AI tools that largely focused on bespoke classification tasks Mello-Thoms and Mello (2023); Alis et al. (2024), Merlin offers a suite of capabilities to support clinicians across various diagnostic, prognostic, and research tasks (Figure 1h). When a patient undergoes an abdominal CT scan prescribed by their provider, Merlin can streamline the clinical workflow. As the radiologist reviews the CT scan, Merlin can generate a preliminary draft of the radiology report and provide a list of common findings, aiding with interpretation efficiency. Additionally, Merlin’s multimodal retrieval capabilities enable content-based image retrieval, allowing finding similar prior studies. This ability can facilitate improved clinical decision-making by providing radiologists and referring providers with relevant patient cases for comparison. Beyond diagnostic tasks, Merlin can assist with opportunistic analysis of the image for diagnosing current or future disorders using its ability of organ segmentation and future disease prediction. Using the abdominal CT image, Merlin can also identify the appropriate ICD-9 and ICD-10 codes, allowing for appropriate diagnostic coding for patients. This automation may reduce the risk of billing errors and claim denials, which can arise from incorrect coding Kusnoor et al. (2020); Weiner (2018).
@@
Recent concurrent research has started to investigate 3D vision language models for radiology, particularly focusing on chest and head CT scans Hamamci et al. (2024a); Yang et al. (2024). However, the extent of experiments and 3D evaluations are limited. These studies lack systematic ablation studies and do not utilize all available clinical data, including EHR. In addition to exploring the use of multiple clinical data types for supervision, we carry out a wide array of experiments and evaluations that contextualize the methodological choices we make in the development of Merlin.
Through the ablation studies, we find that I3D weight initialization Carreira and Zisserman (2017) is helpful for all tasks that we examine with ablation studies. Furthermore, training with both EHR diagnoses and radiology reports is beneficial over training with either alone. The manner of combining these supervision sources is important, with multi-task training using EHR and radiology reports outperforming staged training. Nonetheless, report-only contrastive pretraining accounts for the vast majority of the gains, and the modest additional improvement from incorporating EHR diagnoses may not justify the added complexity in settings where such data are difficult to obtain. Additionally, we find that report splitting, while essential for zero-shot classification, slightly degrades zero-shot retrieval performance. This reinforces that pretraining is most effective when the pretraining data distribution matches the downstream task distribution. Splitting the radiology reports results in text that better aligns with prompts for zero-shot classification, but less closely resembles the radiology report text used for retrieval tasks.
We also demonstrate the advantages of Merlin’s 3D vision-language pretraining strategy over finetuned 2D VLMs, 2D-to-3D lifted VLMs, 3D vision-only models, and recent CT embedding foundation modelsCodella et al. (2024); Yang et al. (2024). Specifically, we find that (i) vision-language pretraining outperforms vision-only pretraining, (ii) finetuning improves the performance of 2D and 2D-to-3D VLMs though they underperform relative to Merlin, (iii) Merlin outperforms recent CT embedding foundation models, and (iv) Merlin outperforms other baselines in both data-scarce and fully supervised settings. Merlin outperformed 3D vision-only SSL and non-SSL methods, suggesting that language alignment provides a stronger self-supervision signal during CLIP pretraining compared to 3D vision-only SSL approaches. Additionally, 2D VLMs exhibited limited capacity to capture volumetric representations required for evaluation tasks, presumably due to their inability to process full 3D volumes. Similarly, 2D-to-3D lifted models failed to effectively encode volumetric information, as their projection layers insufficiently compressed multi-slice data relative to Merlin’s full-volume processing. Furthermore, in the supervised findings-based disease classification task, zero-shot Merlin still outperformed linear probing, mirroring the results observed in the original CLIP experiments Radford et al. (2021a), likely due to its ability to leverage the pretrained alignment between vision and language modalities. This alignment enabled zero-shot inference to generalize well across tasks, even without task-specific finetuning. Overall, our findings suggest that 3D VLM pretraining through Merlin demonstrates trends consistent with other foundation model studies Wornow et al. (2023).
Beyond outperforming alternative architecture baselines, Merlin demonstrates strong generalizability across diverse external datasets. Merlin maintained consistent and strong performance when evaluated on over 44,000 external CT scans spanning multiple sites and anatomies. These results highlight the model’s ability to handle distributional shifts between training and testing cohorts, including differences in patient demographics, acquisition parameters, and reporting practices. Importantly, these findings suggest that Merlin’s pretraining captures features that can generalize across sites, providing a foundational backbone for CT foundation model development and applications.
@@
    Increasing dataset size: Expectedly, our initial data scaling results show that using a larger pretraining dataset improves Merlin’s performance (Figure 2d, Figure 3b, Figure 4e). Our data scaling curves across tasks can serve as a useful prospective reference measure in evaluating adequate sizes of training datasets for a requisite task performance. We note that as we add more data, we may need to update our model architecture, which was optimized at our current data scale.
@@
Beyond the strengths of our study, there are also limitations of our model and our work. First, towards our goal of allowing compute-constrained health systems to train and finetune foundation models, we train and evaluate Merlin using a single GPU. This may limit the generalizability of some of our findings as task performance may increase by simply scaling up compute resources. Second, the limited number of publicly-available baselines for abdominal CT, especially VLM baselines, makes it challenging to understand the relative efficacy of our methods. On the other hand, we believe that Merlin could establish a strong and task-agnostic initial baseline for abdominal CT. Third, further exploration is required to optimize Merlin for the tasks of radiology report generation and 3D segmentation. For radiology report generation, there are numerous parameters that can be tuned, primarily regarding the adapter and LLM. While we devise an initial performant strategy for direct report generation, we leave further optimization of the adapter and LLM to future work. Likewise, adapting Merlin for segmentation requires adding a decoder network. Future work can explore additional architectural parameters and optimization strategies for decoder training.